OpenStudy (anonymous):

If nC10 = nC12 , find 23Cn

hartnn (hartnn):

u know the identity: \(\huge ^nC_r=^nC_{n-r}\)

OpenStudy (anonymous):

Yes

hartnn (hartnn):

so here r=10 n-r=12 n-10=12 n=22 ok?

hartnn (hartnn):

now \(\huge ^{23}C_{22}=^{23}C_{23-22}=^{23}C_{1}=??\)

hartnn (hartnn):

@Viks u know how to find 23C1 ?

OpenStudy (anonymous):

Noo

hartnn (hartnn):

\(\huge ^{n}C_{r}=\frac{n!}{r!(n-r)!}\) so 23C1 will be 23!/(1!*22!)=23*22!/22!=23 ok?

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

If nPr =720 and nCr = 120 Find r